OpenStudy (anonymous):

Subtract and simplify: 2/3x+5 - 1/6x+10

OpenStudy (anonymous):

find the LCM of the two fractions and simplify

OpenStudy (espex):

Is your question \[\frac{2}{3x+5} - \frac{1}{6x+10}\] and @vamgadu is right, find a common denominator for your fractions and then you can manipulate them.

OpenStudy (anonymous):

The following may not be what was intended, however, with no parenthesis, the problem expression is the following:\[\frac{2 x}{3}+5-\frac{x}{6}+10 \]Simplified:\[\frac{x+30}{2} \]